OpenStudy (anonymous):

Please help i dont know how to do this: Verify the identity: cos(x+(pi/2))=-sin(x)

OpenStudy (whpalmer4):

How about using the identity \(\cos(u+v) = \cos u\cos v - \sin u\sin v\) on \(\cos(x+\pi/2)\) and see what develops?

OpenStudy (anonymous):

i replace that with the cos(x+(pi/2)

OpenStudy (anonymous):

?

OpenStudy (whpalmer4):

let u = x, v = pi/2 and expand that identity. it would replace the left hand side of your equation...

OpenStudy (whpalmer4):

Your left hand side is \(\cos (x + \pi/2)\) which is the cosine of a sum. That identity I gave you expands the cosine of a sum.

OpenStudy (anonymous):

i see so i put cos x cos pi/2 - sin x sin pi/2

OpenStudy (anonymous):

then simplify and then ill get the same as the right side?

OpenStudy (whpalmer4):

Well, try it and see!

OpenStudy (whpalmer4):

A lot of these problems you do quite a bit of the learning just by going through the steps and writing them out. Looking at it and saying "oh, yeah, I see" doesn't have the same effect.

OpenStudy (anonymous):

i see but how would i make them alike?

OpenStudy (anonymous):

i need to subsitute cos to make sin

OpenStudy (whpalmer4):

What is the value of \(\cos \pi/2\)? How about \(\sin\pi/2\)?

OpenStudy (anonymous):

cos(pi/2)= 0

OpenStudy (whpalmer4):

Yes. What about sin?

OpenStudy (anonymous):

1

OpenStudy (whpalmer4):

Okay, so what does the left half become after you substitute in those values?

OpenStudy (anonymous):

cos(x)-sin(x)+1

OpenStudy (anonymous):

so now we have to get rid of the cos(x) and the +1

OpenStudy (whpalmer4):

No...you have\[\cos(x+\pi/2) = -\sin x\] We substitute the angle sum identity for cosines on the left, with u = x, and v = pi/2:\[\cos x\cos \pi/2 - \sin x\sin \pi/2 = -\sin x\]\[\cos \pi/2 = 0\]\[\sin \pi/2 = 1\]Substitute those values into the equation \[0*\cos x - 1 *\sin x = -\sin x\]\[-\sin x = -\sin x\]